2, For the following periodic series, f(t) = t^2, ( -1<t<1), period = 2

(a) graph this periodic function, range is ( -8<t<8), and find a key point

(b)Even function or odd function?

(c ) find a Fourier series
Answer #1

Solna Giyen, flt) 2 + a f(a)=Y fitry o - 3 -2 - + 4 s 3 6 7 Graph of fitl=t2 with period 2 ico f(t) = f(t+2) yt (b) f(+226? aSo 902 dj- 4C4). Smo dt = 2 Xo ds 20 and ana Ť S fck). Son mpit dat 13. Sine is cell function) s t2Sm ntt at a f (even) xfltl2 0+ Eox Sm 20 + Q cin n21 na 2012 Cosnat fitla Enn 3 5 nal Cosnit (hr)? Angwen CScanned with CamScanner